What is the ksp expression for PBI2 when 1.5 M Cal2 is added?
1
Expert's answer
2020-10-22T08:39:08-0400

the ksp expression for PBI2 

PbI2"\iff" Pb2+ + 2I-

x 2x

Ksp= [Pb]2+ [I-]2

Value of Ksp is 1.4 × 10-8

1.4×10-8= 4x3

X3=3.5×10-9

X= 1.1518×10-3

So the concentration of I- is

3.036×10-3

Let the PbI2 added in 1.5 M NaI

[Pb2+] = sM

[I-] = (2s + 1.5)M

Ksp for PbI2 is

Ksp= [sM] [2s+1.5]M